# LoGeR: Long-Context Geometric Reconstruction with Hybrid Memory

### 03 Mar 2026 v1

@@ 1 Introduction
Furthermore, to evaluate these mechanisms at an unprecedented scale, we introduce a long-context reconstruction benchmark derived from the VBR dataset (Brizi et al., 2024), spanning up to 19k frames and 11.5 km in trajectory. Remarkably, our architectural design enables the model to be trained on sequences of 128 frames, generalize seamlessly up to 1k frames during inference, and scale effectively up to 19k frames with periodic state resets and an optional feedforward pose-alignment. As shown in Fig. 1, LoGeR effectively breaks both the context and data walls. It substantially outperforms prior state-of-the-art feedforward methods on KITTI, reducing the Absolute Trajectory Error (ATE) from 72.86 to 18.65, and achieves a 30.8% relative improvement on the extremely long sequences of our VBR benchmark.
@@ 2 Related Work
Long sequence reconstruction. Recent feedforward approaches tackle long sequence reconstruction using external spatial memory (Chen et al., 2025; Wang & Agapito, 2025; Wu et al., 2025; Lan et al., 2025), RNN-style persistent state (Wang et al., 2025b), or efficiency-oriented VGGT variants with sparse (Shen et al., 2026; Yuan et al., 2026) or causal (Zhuo et al., 2026) attention. However, their lack of demonstrated results on exceptionally long sequences (e.g., minute-long) leaves their scalability in sequence lengths unverified. Concurrently, TTT3R (Chen et al., 2026) introduces a confidence-based update for single-frame streaming (Wang et al., 2025b). However, this frame-wise linear approach lacks the expressivity to capture complex temporal context or leverage the powerful multi-frame reasoning of bidirectional backbones (Wang et al., 2025a, 2026).
@@ 4.2 Learning Objectives
$\mathcal{L}_{\text{global}}$ $\!=\!\frac{1}{N|\Omega|}\sum_{i=1}^{N}\sum_{p\in\Omega}\big\|\,\Pi(\hat{\mathbf{T}}_{i},\,\hat{\mathbf{x}}_{i,p})\!-\!\Pi(\mathbf{T}_{i},\,\mathbf{x}_{i,p})\,\big\|_{1},$ (10)
@@
Figure 4: Quantitative results on our proposed VBR (Brizi et al., 2024) evaluation showing results on very long sequences spanning from 1,000 to 19,000 frames. Our methods achieve 30.8% more accurate results than prior methods.
@@ Acknowledgment
We would especially like to thank Noah Snavely for helpful feedback throughout the project. We thank Tianyuan Zhang, Songlin Yang, Youming Deng, Haiwen Feng, Qianqian Wang, and Yifei Zhang for helpful discussions. We thank Alfred Piccioni for help with the training infrastructure. We thank Xingyu Chen for providing details on the evaluation. We thank Angjoo Kanazawa, Tyler Bonnen, and Jiahui Lei for helpful feedback on the manuscript.
@@ A.4 More Inference Details
Opt. VGGT-SLAM (Maggio et al., 2025) 10.05 9.67 8.47 8.15 7.55 5.80 11.14 8.69
VGGT-Long w/o LC (Deng et al., 2025) 9.03 10.89 9.91 7.31 6.85 5.55 13.05 8.94
VGGT-Long (Deng et al., 2025) 6.29 10.89 9.91 7.31 7.09 5.55 13.12 8.60
Feedforward InfiniteVGGT (Yuan et al., 2026) 9.16 11.12 10.00 8.41 7.50 5.62 13.70 9.36
CUT3R (Wang et al., 2025b) 9.09 6.50 6.57 6.83 6.68 5.35 13.26 7.75
TTT3R (Chen et al., 2026) 8.69 7.71 7.52 5.82 6.11 4.30 13.18 7.62
Pi3-Chunk (Proposed Baseline) 8.78 8.86 8.11 6.48 6.69 4.88 10.57 7.77
LoGeR (Ours) 4.88 4.13 5.18 5.03 4.55 5.75 8.30 5.40
LoGeR* (Ours) 6.92 4.60 5.59 3.18 4.67 5.48 6.44 5.27

### 27 Apr 2026 v2

@@ 1 Introduction
Furthermore, to evaluate these mechanisms at an unprecedented scale, we introduce a long-context reconstruction benchmark derived from the VBR dataset (Brizi et al., 2024), spanning up to 19k frames and 11.5 km in trajectory. Remarkably, our architectural design enables the model to be trained on sequences of 128 frames, generalize seamlessly up to 1k frames during inference, and scale effectively up to 19k frames with periodic state resets and an optional feedforward pose-alignment. As shown in Fig. 1, LoGeR effectively breaks both the context and data walls. It substantially outperforms prior state-of-the-art feedforward methods on KITTI, reducing the Absolute Trajectory Error (ATE) from 72.86 to 18.65, and achieves a 55.2% relative improvement on the extremely long sequences of our VBR benchmark.
@@ 2 Related Work
Long sequence reconstruction. Recent feedforward approaches tackle long sequence reconstruction using external spatial memory (Chen et al., 2025; Wang & Agapito, 2025; Wu et al., 2025; Lan et al., 2025), RNN-style persistent state (Wang et al., 2025b), or efficiency-oriented VGGT variants with sparse (Shen et al., 2026; Yuan et al., 2026) or causal (Zhuo et al., 2026; Cheng et al., 2026) attention. However, their lack of demonstrated results on exceptionally long sequences (e.g., minute-long) leaves their scalability in sequence lengths unverified. Concurrently, TTT3R (Chen et al., 2026) introduces a confidence-based update for single-frame streaming. However, this frame-wise linear approach lacks the expressivity to capture complex temporal context or leverage the powerful multi-frame reasoning of bidirectional backbones (Wang et al., 2025a, 2026).
@@ 4.2 Learning Objectives
$\mathcal{L}_{\text{global}}$ $\!=\!\frac{1}{N|\Omega|}\sum_{i=1}^{N}\sum_{p\in\Omega}\big\|\,s^{\ast}\Pi(\hat{\mathbf{T}}_{i},\,\hat{\mathbf{x}}_{i,p})\!-\!\Pi(\mathbf{T}_{i},\,\mathbf{x}_{i,p})\,\big\|_{1},$ (10)
@@
Figure 4: Quantitative results on our proposed VBR (Brizi et al., 2024) evaluation showing results on very long sequences spanning from 1,000 to 19,000 frames. Our methods achieve 55.2% more accurate results than prior methods.
@@ Acknowledgment
We would especially like to thank Noah Snavely for helpful feedback throughout the project. We thank Tianyuan Zhang, Songlin Yang, Youming Deng, Haiwen Feng, Qianqian Wang, and Yifei Zhang for helpful discussions. We thank Alfred Piccioni for the help with the training infrastructure. We thank Xingyu Chen for providing details on the evaluation. We thank Zihan Zhu for his feedback on the evaluation of the VBR benchmark. We thank Angjoo Kanazawa, Tyler Bonnen, and Jiahui Lei for helpful feedback on the manuscript.
@@ A.4 More Inference Details
Opt. VGGT-SLAM (Maggio et al., 2025) 102.92 110.77 89.69 72.98 62.67 35.62 144.02 88.38
VGGT-Long w/o LC (Deng et al., 2025) 86.46 132.66 115.65 64.40 57.39 33.66 186.13 96.62
VGGT-Long (Deng et al., 2025) 45.73 132.68 115.65 64.40 58.49 33.66 187.97 91.23
Feedforward InfiniteVGGT (Yuan et al., 2026) 90.02 136.83 116.69 77.72 63.47 33.91 202.11 102.96
CUT3R (Wang et al., 2025b) 88.16 46.39 48.00 52.87 47.98 30.18 187.58 71.59
TTT3R (Chen et al., 2026) 81.63 68.36 65.67 35.56 39.03 19.84 185.59 70.81
Pi3-Chunk (Proposed Baseline) 81.63 86.54 71.49 49.85 54.27 28.07 118.81 70.09
LoGeR (Ours) 25.60 19.53 28.90 28.11 23.04 35.01 92.94 36.16
LoGeR* (Ours) 49.38 22.44 34.90 11.08 27.05 32.88 44.51 31.75